Test file for Victor’s outline macros

         Victor Eijkhout, victor@eijkhout.net

May 20, 1999

Outline test

First section

This is just an example, remember? • A point. But it should be sufficient.

• Another point. Even if you just want to copy from this file. something: is useful, another thing: is not.

Another section, another page

• This one has this.

• And that..

This could be an included chapter

by another author.

• This point is level 1 by the author, 2 by inclusion. Yadah yadah yadah.

• Another level 1 point made level 2. Note that this time there is no page break.

Another chapter, by another author

More text.

Conclusions

So, do you like it?